OpenStudy (anonymous):

The slope of the line below is -4. Write the equation of the line in point-slope form, coordinates (1,2)

OpenStudy (anonymous):

(x1,y1) = (1,2) m=-4 point slope form is given by (y-y1)=m(x-x1)

OpenStudy (anonymous):

Ok so we have y-y1=m(x-x1) What's the M in this equation?

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

k good and what's the x1 & y1?

OpenStudy (anonymous):

y is 1 x is 2

OpenStudy (anonymous):

close but actually other way around :) (x1,y1)= (1,2) = x1=1 & y1=2

OpenStudy (anonymous):

Ok so now that we have that, can u try putting the numbers in the equation where they belong? :)

OpenStudy (anonymous):

o ya lookd at it wrong

OpenStudy (anonymous):

y-2=-4(x-1) is what i came up with

OpenStudy (anonymous):

yep perfect! :) & yeah I thought u looked at it wrong or something lol, i knew u knew that :) Great job :D

OpenStudy (anonymous):

it is - right not +

OpenStudy (anonymous):

U mean -4 instead of 4? Yeah because in the problem they said slope is -4, we can't take off that negative sign that was already on it :)

OpenStudy (anonymous):

alright il post in question in live feed